# Interesting Scientific Idea Generation Using Knowledge Graphs and LLMs: Evaluations with 100 Research Group Leaders

### v2

@@ Introduction
Figure 1: SciMuse suggests research ideas or collaborations using a knowledge graph and GPT-4. (a), Knowledge graph generation. Nodes represent scientific concepts extracted from 2.44 million paper titles and abstracts using the RAKE algorithm [22], further refined with custom NLP techniques, manual review, GPT, and Wikipedia (to restore mistakenly removed concepts), resulting in a final list of 123,128 concepts. Edges are formed when two concepts co-occur in titles or abstracts of over 58 million papers from OpenAlex [23], augmented with citation data as a proxy for impact. A mini-knowledge graph illustrates the connections for two example papers [24, 25]. (b), AI-generated research collaborations. We extract concepts from the publications of Researchers A and B, refine them using GPT-4, and identify relevant sub-networks in the knowledge graph. GPT-4 then uses these concept pairs, along with the researchers’ research information, to generate personalized research ideas or collaboration projects.
@@ Results
Figure 2: Large-scale human evaluation within the Max Planck Society. (a)-(b), The map of Germany, based on the GISCO statistical unit dataset from Eurostat [28], shows the locations of the Max Planck Institutes and the participating group leaders. A total of 4,451 personalized AI-generated research suggestions were evaluated by 110 research group leaders. Each suggestion represents a potential collaboration between the evaluating researcher (Researcher A) and another researcher (Researcher B) from the Max Planck Society, visualized as bi-colored edges (orange for Researcher A, green for Researcher B). A purple circle indicates collaborations within the same institute, and edge transparency reflects the number of evaluated suggestions. Blue dots denote natural sciences (nat) and red dots represent social sciences (soc). (c): Distribution of interest ratings on a scale from 1 (‘not interesting’) to 5 (‘very interesting’), with 394 suggestions rated as very interesting and 713 rated 4. Ratings are further categorized by whether the collaborations are within or across institutes, and by research field affiliation in either the natural or social sciences.
The knowledge graph, depicted in Fig. 1(a), consists of vertices, representing scientific concepts, and edges are drawn when two concepts jointly appear in a title or abstract of a scientific paper. The concept list is generated from the titles and abstracts of approximately 2.44 million papers from arXiv, bioRxiv, ChemRxiv, and medRxiv, with a data cutoff in February 2023. Rapid Automatic Key-word Extraction (RAKE) algorithm based on statistical text analysis is used to extract candidate concepts [22]. Those candidates are further refined using GPT, Wikipedia, and human annotators, resulting in 123,128 concepts in the natural and social sciences. We then use more than 58 million scientific papers from the open-source database OpenAlex [23] to create edges. These edges contain information about the co-occurrence of concepts in scientific papers (in titles and abstracts) and their subsequent citation rates. This new knowledge graph representation was recently introduced in [12] to predict the impact of future research topics. As a result, we have an evolving knowledge graph that captures part of the evolution of science from 1665 (a text by Robert Hooke on the observation of a great spot on Jupiter [29]) to April 2023. Details of the knowledge graph generation are shown in Fig. 1(a) and the Appendix.
@@
With the researchers’ subgraphs, we then generate a prompt for GPT-4 to create a research project (details in the Appendix). In the prompt, we provide the titles of up to seven papers from each researcher and ask GPT-4 to create a research project based on two selected scientific concepts. These concepts are chosen in one of three ways: (1) a randomly selected concept pair, (2) the concept pair with the highest predicted impact, or (3) no specific concept pair, relying only on paper titles. We predict the impact (in terms of expected future citations) by adapting the computational methods from [12], applying them to a different and much larger knowledge graph. While the third method does not directly use the knowledge graph features, it serves as a valuable sanity check for our approach (see Appendix). The prompt incorporates self-reflection techniques [30], where GPT-4 generates three ideas, iteratively refines them twice, and then selects the most suitable one as the final result.
Figure 3: Analysis of interest levels versus knowledge graph features. We analyzed how eight features of the knowledge graph correlate with researchers’ interest levels. After normalizing these features using z-scores, we arranged them from lowest to highest and divided the data into 50 equal groups. For each group, we plotted the average feature value (x-axis) against the average interest level (y-axis) with standard deviations, to identify trends. (a) and (b) show node features, (c)–(e) show node citation metrics, (f) shows an edge feature, (g) an edge citation metric, and (h) represents semantic distance between researchers’ sub-networks (higher values indicate that the researchers’ scientific fields are further apart). Data points include all 2,996 responses (blue), the top 50% of concept pairs by predicted impact (green), and the top 25% (red).
@@
Figure 4: Predicting Scientific Interest. We use two distinct methods to predict interest levels: (1) a supervised neural network trained on human evaluations using only knowledge graph data (not the text of the actual suggestion), and (2) GPT in a zero-shot setting, ranking suggestions without getting any feedback from human evaluations. Both methods classify suggestions as highly interesting (ratings of 4 or 5) or not (below 4). The neural network uses 25 knowledge graph features and employs Monte Carlo cross-validation for accuracy estimation. For GPT, we conduct pairwise comparisons using personalized research details and rank suggestions through an ELO-based tournment system. (a), The ROC curve shows prediction accuracy of 63.9% for the neural network and 67.3% for GPT-4o. (b), The precision for top-N suggestions is significantly higher than random selection, with the top-1 precision reaching 70% for the neural network (51.0% for GPT-4o and 52.9% for GPT-3.5) and top-5 precision at 60.4% (46.7% for GPT-4o, 43.7% for GPT-3.5). (c), The probability of having at least one high-interest suggestion among the top N recommendations is significantly higher for the supervised neural network compared to random selection. Practically, evaluation data from experienced researchers may not always be available, thus it is very encouraging that LLMs, even without human evaluation, can rank suggestions effectively such that the highest interesting ones appear first.
@@
For the supervised neural network, we used knowledge graph features to predict whether a research proposal would receive a high rating (4 or 5) or below 4. Given the limited training data – each of the 2,996 data point representing a research group leader’s evaluation of a proposal’s interest level – we employed a low-data machine learning approach with a small neural network (25 high-performing input features, 50 neurons in a single hidden layer, and one output neuron). Dropout was used for training [31], and Monte Carlo cross-validation [32] (also known as repeated random sub-sampling validation) was applied to ensure robust evaluation and maximize the utility of our limited data (see the Appendix).
@@ Acknowledgements
The authors wholeheartedly thank all the researchers who spent the time participating in our study. The authors also thank the organizers of OpenAlex, arXiv, bioRxiv, and medRxiv for making scientific resources freely accessible. X.G. acknowledges support from the Alexander von Humboldt Foundation.
@@
Over 100 highly experienced researchers, spanning fields from the natural sciences to the humanities, participated in evaluating the personalized research ideas. Table. 1 summarizes the researchers’ publication and citation statistics as of January 1, 2024, when the evaluations were conducted. On average, the researchers had published 59 papers and received over 3,750 citations.
Table 1: Summary statistics of researchers’ publications and citations.
@@
The prompt used to suggest research ideas based on concept pairs from the knowledge graph is described in the following Box:
@@
Fig. 5 presents the interest-level distributions for research suggestions generated using three different methods. The interest levels are notably similar between suggestions generated with and without concepts from the knowledge graph. This similarity enables us to analyze the correlations between knowledge graph properties and interest levels, and to use these properties for predicting the interest level of generated research proposals.
Figure 5: Interest levels across different generation methods. Research ideas are generated using three methods: (1) no concepts provided by the knowledge graph, (2) random concepts from the researchers’ subnetwork, and (3) predicted high-impact concept pairs from the researchers’ subnetwork. The figures displays: (a) overall interest levels (numbers within bars show the number of responses for that evaluation), (b) interest levels for ideas without using concepts from the knowledge graph, (c) interest levels with random concept pairs, and (d) interest levels using high-impact concept pairs (predicted by adapting the computational methods from [12], and applying them to a different and much larger knowledge graph).
.8 Predicting high interest from knowledge graph features
@@
The input to the neural networks consists of network-theoretical features extracted from the knowledge graph. For each concept pair in a research project, we compute 144 features. The first 141 features are derived from those used to predict the future impact of concept pairs, as described in [12]. These features include node properties (e.g., node degree and PageRank [43]) and edge properties (e.g., Simpson similarity and Sørensen–Dice coefficient [44]). Several features also account for impact information, such as recent citation counts. The remaining three features include the predicted impact and two distance metrics between the researchers’ subgraphs (Fig. 1(b)). The first distance metric measures the distance based solely on the concepts present in Researcher A and Researcher B’s concept lists. In contrast, the second metric takes into account the entire neighborhood of these subgraphs by calculating semantic distances between all neighboring concepts and the concepts from the subgraphs. These features serve as the input to the neural network for predicting whether a proposal will achieve a high interest rating.
Given the small dataset size (2,996 answers with properties from the knowledge graph), we use a data-efficient learning method – a small neural network with dropout. The input layer consists of the 25 best-performing features (see Table. 2), selected from the total 144 by independently analyzing the feature importance of each and choosing the top 25. The neural network has one hidden layer with 50 neurons and a single output neuron. Mean square error is used as the loss function.
@@
The neural network performance is not specifically sensitive to hyperparameter choices, thus we refrained from hyperparameter optimization, and instead used a reasonable defaults: learning rate=0.003, dropout=20%, weight decay=0.0007, training dataset=75%, validation dataset=15%, test dataset=10%. We also experimented with other data-efficient learning methods, such as decision trees, but found that they did not significantly outperform the neural network.
Table 2: Top-25 Best-Performing Features of Each Concept Pair ($c_{A}$, $c_{B}$) Input to the Neural Network
@@ 25 Total citations for $c_{A}$ from its first publication until the year 2022
.9 Zero-shot ranking of research suggestions by GPT
We ranked 2,996 research suggestions – previously evaluated by human experts – using GPT-3.5, GPT-4o, and GPT-4o-mini. For each pair of randomly selected suggestions, we asked the LLMs to rank which one was more interesting, considering the personalized research interests of the evaluating human expert. This pairwise comparison was repeated between 22,000 and 45,000 times (for GPT4o and GPT4o-mini, respectively). We treated this task as a tournament where all 2,996 suggestions compete pairwise against each other. Using the ELO ranking system, each suggestion started with an initial ELO score of 1400. Each comparison by GPT updated the ELO rankings based on the outcome, producing a final sorted list of suggestions from highest to lowest ELO score. We evaluated the ranking quality by calculating the AUC to determine how well the ranked list aligns with the human-expert evaluations of interest levels (Fig.6). The prompt we use for the zero-Shot ranking of research suggestions is shown in the Box.
Figure 6: Zero-Shot ranking of research suggestions by LLMs. The research suggestions are generated using the knowledge graph together with GPT4. They are then ranked using GPT4o, GPT4o-mini and GPT3.5, without feedback from the human evaluation. The human evaluation is used to compute the final quality of the ranking. the ranking is performed in a pair-wise choice where we ask the LLM to select the more interesting one given the research background of the researchers. One match is one pairwise selection. The LLMs perform 10,000 of these pairwise selections, which allows us to compute ELO scores for each generated research idea.

### v3

@@ Introduction
Fig. 1: SciMuse suggests research ideas or collaborations using a knowledge graph and GPT-4. (a), Knowledge graph generation. Nodes represent scientific concepts extracted from 2.44 million paper titles and abstracts using the RAKE algorithm [22], further refined with custom NLP techniques, manual review, GPT, and Wikipedia (to restore mistakenly removed concepts), resulting in a final list of 123,128 concepts. Edges are formed when two concepts co-occur in titles or abstracts of over 58 million papers from OpenAlex [23], augmented with citation data as a proxy for impact. A mini-knowledge graph illustrates the connections for two example papers [24, 25]. (b), AI-generated research collaborations. We extract concepts from the publications of Researchers A and B, refine them using GPT-4, and identify relevant sub-networks in the knowledge graph. GPT-4 then uses these concept pairs, along with the researchers’ research information, to generate personalized research ideas or collaboration projects.
@@ Results
Fig. 2: Large-scale human evaluation within the Max Planck Society. (a)-(b), The map of Germany, based on the GISCO statistical unit dataset from Eurostat [28], shows the locations of the Max Planck Institutes and the participating group leaders. A total of 4,451 personalized AI-generated research suggestions were evaluated by 110 research group leaders. Each suggestion represents a potential collaboration between the evaluating researcher (Researcher A) and another researcher (Researcher B) from the Max Planck Society, visualized as bi-colored edges (orange for Researcher A, green for Researcher B). A purple circle indicates collaborations within the same institute, and edge transparency reflects the number of evaluated suggestions. Blue dots denote natural sciences (nat) and red dots represent social sciences (soc). (c): Distribution of interest ratings on a scale from 1 (‘not interesting’) to 5 (‘very interesting’), with 394 suggestions rated as very interesting and 713 rated 4. Ratings are further categorized by whether the collaborations are within or across institutes, and by research field affiliation in either the natural or social sciences.
The knowledge graph, depicted in Fig. 1(a), consists of vertices, representing scientific concepts, and edges are drawn when two concepts jointly appear in a title or abstract of a scientific paper. The concept list is generated from the titles and abstracts of approximately 2.44 million papers from arXiv, bioRxiv, ChemRxiv, and medRxiv, with a data cutoff in February 2023. Rapid Automatic Key-word Extraction (RAKE) algorithm based on statistical text analysis is used to extract candidate concepts [22]. Those candidates are further refined using GPT, Wikipedia, and human annotators, resulting in 123,128 concepts in the natural and social sciences. We then use more than 58 million scientific papers from the open-source database OpenAlex [23] to create edges. These edges contain information about the co-occurrence of concepts in scientific papers (in titles and abstracts) and their subsequent citation rates. This new knowledge graph representation was recently introduced in [12] to predict the impact of future research topics. As a result, we have an evolving knowledge graph that captures part of the evolution of science from 1665 (a text by Robert Hooke on the observation of a great spot on Jupiter [29]) to April 2023. Details of the knowledge graph generation are shown in Fig. 1(a) and Supplementary Information.
@@
With the researchers’ subgraphs, we then generate a prompt for GPT-4 to create a research project (details in the Supplementary Information). In the prompt, we provide the titles of up to seven papers from each researcher and ask GPT-4 to create a research project based on two selected scientific concepts. These concepts are chosen in one of three ways: (1) a randomly selected concept pair, (2) the concept pair with the highest predicted impact, or (3) no specific concept pair, relying only on paper titles. We predict the impact (in terms of expected future citations) by adapting the computational methods from [12], applying them to a different and much larger knowledge graph. While the third method does not directly use the knowledge graph features, it serves as a valuable sanity check for our approach (see Supplementary Information). The prompt incorporates self-reflection techniques [30], where GPT-4 generates three ideas, iteratively refines them twice, and then selects the most suitable one as the final result.
Fig. 3: Analysis of interest levels versus knowledge graph features. We analyzed how eight features of the knowledge graph correlate with researchers’ interest levels. After normalizing these features using z-scores, we arranged them from lowest to highest and divided the data into 50 equal groups. For each group, we plotted the average feature value (x-axis) against the average interest level (y-axis) with standard deviations, to identify trends. (a) and (b) show node features, (c)–(e) show node citation metrics, (f) shows an edge feature, (g) an edge citation metric, and (h) represents semantic distance between researchers’ sub-networks (higher values indicate that the researchers’ scientific fields are further apart). Data points include all 2,996 responses (blue), the top 50% of concept pairs by predicted impact (green), and the top 25% (red), using the neural-network based impact prediction presented in [12].
@@
Fig. 4: Predicting Scientific Interest. We use two distinct methods to predict interest levels: (1) a supervised neural network trained on human evaluations using only knowledge graph data (not the text of the actual suggestion), and (2) GPT in a zero-shot setting, ranking suggestions without getting any feedback from human evaluations. Both methods classify suggestions as highly interesting (ratings of 4 or 5) or not (below 4). The neural network uses 25 knowledge graph features and employs Monte Carlo cross-validation for accuracy estimation. For GPT, we conduct pairwise comparisons using personalized research details and rank suggestions through an ELO-based tournment system. (a), The ROC curve shows prediction accuracy of 64.5% for the neural network and 67.3% for GPT-4o. (b), The precision for top-N suggestions is significantly higher than random selection, with the top-1 precision reaching 70% for the neural network (51.0% for GPT-4o and 52.9% for GPT-3.5) and top-5 precision at 60.4% (46.7% for GPT-4o, 43.7% for GPT-3.5). (c), The probability of having at least one high-interest suggestion among the top N recommendations is significantly higher for the supervised neural network compared to random selection. Practically, evaluation data from experienced researchers may not always be available, thus it is very encouraging that LLMs, even without human evaluation, can rank suggestions effectively such that the highest interesting ones appear first.
@@
For the supervised neural network, we used knowledge graph features to predict whether a research proposal would receive a high rating (4 or 5) or below 4. Given the limited training data – each of the 2,996 data point representing a research group leader’s evaluation of a proposal’s interest level – we employed a low-data machine learning approach with a small neural network (25 high-performing input features, 50 neurons in a single hidden layer, and one output neuron). Dropout was used for training [31], and Monte Carlo cross-validation [32] (which is also known as repeated random sub-sampling validation) was applied to ensure robust evaluation and maximize the utility of our limited data (see the Supplementary Information). Decision trees [33] were not able to outperform the quality of neural networks (see Supplementary Information).
@@ Acknowledgements
The authors wholeheartedly thank all the researchers who spent the time participating in our study. The authors also thank the organizers of OpenAlex, arXiv, bioRxiv, medRxiv, and chemRxiv for making scientific resources freely accessible. X.G. acknowledges support from the Alexander von Humboldt Foundation.
@@
Over 100 highly experienced researchers, spanning fields from the natural sciences to the humanities, participated in evaluating the personalized research ideas. Table. S1 summarizes the researchers’ publication and citation statistics as of January 1, 2024, when the evaluations were conducted. On average, the researchers had published 59 papers and received over 3,750 citations.
Table S1: Summary statistics of researchers’ publications and citations.
@@
The prompt used to suggest research ideas based on concept pairs from the knowledge graph is described as the follows:
@@
Fig. S1 presents the interest-level distributions for research suggestions generated using three different methods. The interest levels are notably similar between suggestions generated with and without concepts from the knowledge graph. This similarity enables us to analyze the correlations between knowledge graph properties and interest levels, and to use these properties for predicting the interest level of generated research proposals.
Fig. S1: Interest levels across different generation methods. Research ideas are generated using three methods: (1) no concepts provided by the knowledge graph, (2) random concepts from the researchers’ subnetwork, and (3) predicted high-impact concept pairs from the researchers’ subnetwork. The figures displays: (a) overall interest levels (numbers within bars show the number of responses for that evaluation), (b) interest levels for ideas without using concepts from the knowledge graph, (c) interest levels with random concept pairs, and (d) interest levels using high-impact concept pairs (predicted by adapting the computational methods from [12], and applying them to a different and much larger knowledge graph).
Fig. S2: Choice of alternative hyper-parameters for training of neural network. We analyse the prediction of interest level quality (in terms of AUC) with different parameters of the neural network, such as different number of features, different number of layers and neurons, learning rate and drop-out rate. We see that the final results are robust under variations of the hyper-parameters.
.8 Predicting high interest from knowledge graph features with neural networks
@@
The input to the neural networks (using PyTorch [44]) consists of network-theoretical features extracted from the knowledge graph. For each concept pair in a research project, we compute 144 features. The first 141 features are derived from those used to predict the future impact of concept pairs, as described in [12]. These features include node properties (e.g., node degree and PageRank [45]) and edge properties (e.g., Simpson similarity and Sørensen–Dice coefficient [46]). Several features also account for impact information, such as recent citation counts. The remaining three features include the predicted impact and two distance metrics between the researchers’ subgraphs (Fig. 1(b)). The first distance metric measures the distance based solely on the concepts present in Researcher A and Researcher B’s concept lists. In contrast, the second metric takes into account the entire neighborhood of these subgraphs by calculating semantic distances between all neighboring concepts and the concepts from the subgraphs. These features serve as the input to the neural network for predicting whether a proposal will achieve a high interest rating.
Given the small dataset size (2,996 answers with properties from the knowledge graph), we use a data-efficient learning method – a small neural network with dropout. The input layer consists of the 25 best-performing features (see Table. S2), selected from the total 144 by independently analyzing the feature importance of each and choosing the top 25. The neural network has one hidden layer with 50 neurons and a single output neuron. Mean square error is used as the loss function.
@@
The neural network performance is not specifically sensitive to hyperparameter choices, thus we refrained from hyperparameter optimization, and instead used a reasonable defaults: learning rate=0.003, dropout=20%, weight decay=0.0007, training dataset=75%, validation dataset=15%, test dataset=10%. In Fig.S2, we investigate alternative hyper-parameters of the training process, and find that the results are robust under variations of the hyper-parameters.
Fig. S3: Choice of hyper-parameters for training of decision tree. The model is trained using Monte Carlo cross-validation until a statistical uncertainty of $\sigma$=0.001 is reached. We find that no setting of number of features, maximum depth and minimal sample leaf can reach the performance of the data-efficient neural network.
.9 Predicting high interest from knowledge graph features with decision trees
We experimented with other data-efficient learning methods, specifically with decision trees [33] using[47]. However, decision trees did not outperform the neural network predictions, as can be seen in Fig.S3. These values confirm that neural networks are advantageous for the task of ranking research ideas by their interest value in a supervised way, which can also be confirmed in Fig.S5.
Fig. S4: Zero-Shot ranking of research suggestions by LLMs. The research suggestions are generated using the knowledge graph together with GPT4. They are then ranked using GPT4o, GPT4o-mini and GPT3.5, without feedback from the human evaluation. The human evaluation is used to compute the final quality of the ranking. the ranking is performed in a pair-wise choice where we ask the LLM to select the more interesting one given the research background of the researchers. One match is one pairwise selection. The LLMs perform 10,000 of these pairwise selections, which allows us to compute ELO scores for each generated research idea.
.10 Zero-shot ranking of research suggestions by GPT
We ranked 2,996 research suggestions – previously evaluated by human experts – using GPT-3.5, GPT-4o, and GPT-4o-mini. For each pair of randomly selected suggestions, we asked the LLMs to rank which one was more interesting, considering the personalized research interests of the evaluating human expert. This pairwise comparison was repeated between 22,000 and 45,000 times (for GPT4o and GPT4o-mini, respectively). We treated this task as a tournament where all 2,996 suggestions compete pairwise against each other. Using the ELO ranking system, each suggestion started with an initial ELO score of 1400. Each comparison by GPT updated the ELO rankings based on the outcome, producing a final sorted list of suggestions from highest to lowest ELO score. We evaluated the ranking quality by calculating the AUC to determine how well the ranked list aligns with the human-expert evaluations of interest levels (Fig.S4).
.11 Prediction of Interest with different methods
We show the full data of all five methods (supervised training with neural networks and decision trees, as well as unsupervised zero-shot prediction with GPT3.5, GPT4o and GPT4o-mini), with their corresponding AUC, top-N precision and top-N success probability, in Fig.S5. We see that the neural network outperforms decision trees when trained in a supervised way, and that GPT4o is better than the other tested models, when the ranking is performed in a zero-shot manner without giving any information about the evaluations of humans.
.12 Prompt engineering
We have explored manual and automated improvements of the prompts, both for the research question design and the zero-shot prediction. Specifically, we attempted to improve the prompts for the idea generation using GPT-4o. While the prompts were more structured, a small-scale evaluation did not show any improvement in terms of more interesting results.
.13 GPT-4o and GPT-o1 for idea generation
We conducted two small-scale tests where GPT-4 and GPT-4o generated ideas using the exact same settings described above.
In the first test, three research group leaders evaluated 180 pairs of questions (one generated by GPT-4 and the other by GPT-4o using the same prompt). They found 31.1% of GPT-4 answers to be more interesting, while 60.56% favored GPT-4o answers (8.3% were draw). In the second test, a research group leader evaluated 11 pairs of questions (one generated by GPT-4o and the other by GPT-o1 with the same prompt), and all 11 ideas generated by GPT-o1 were ranked as more interesting.
These additional small-scale tests suggest that improved models can enhance idea generation, thus directly improving the results of SciMuse.
Fig. S5: Interest predictions with five methods. We reproduce Fig.4 (main text), and add results from a supervised decision tree training, and an unsupervised GPT4o-mini.
Table S2: Top-25 Best-Performing Features of Each Concept Pair ($c_{A}$, $c_{B}$) Input to the Neural Network